.Air at 1 bar, 27oC and a rate of 0.4 kg/s enters a 10 cm diameter duct where it 

is heated at constant pressure to reach 727oC at exit. Calculate outlet velocity 

and deduce heat and work exchanged per second.


1
Expert's answer
2021-07-05T16:05:34-0400

Gives


"P_1=1atm,T_1=300K,r_1=5cm,\frac{m}{dt}=0.4kg/sec,T_2=727°c=1000K"

We know that

"Av=\frac{dm}{dt}"

Put value


"A=\pi r^2=3.14\times0.05^2=7.85\times10^{-3}m^2"

"v=\frac{\frac{dm}{dt}}{A}"

"v=\frac{0.4}{7.85\times10^{-3}}=50.95m/sec"